OpenStudy (anonymous):

85% of 640 programs is how many?

OpenStudy (anonymous):

544

OpenStudy (anonymous):

85 % of 640 implies 85/100 * 640 = 0.85 * 640 = 544

OpenStudy (anonymous):

whenever u have percentage that means ÷ 100

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

85% * 640 = 544